jamil is 3/8 of Ajmal's age, if Jamil is 15 how old is Ajmal
Answer by mananth(16946)

Jamil is 15 years
Jamil is 3/8 of Ajmal's age
let ajmals age be x
Jamil is 3/8 *x
3/8*x = 15
multiply by 8
3/8 * x * 8 = 15*8
3x=120
/3
x=120/3
x=40 years Ajmal's age
...
m.ananth@hotmail.ca